OpenStudy (anonymous):

Rental car agency charges $190 per week plus 15 cent per mile to rent a car. How many miles can you travel in one week for $265?

OpenStudy (lilg132):

265 - 190 = $75 75 / 0.15 = 500 miles remember to click good answer

OpenStudy (anonymous):

500 miles.

OpenStudy (karatechopper):

190+0.15n=265 190-190+0.15n=265-190 0.15n=75 0.15/0.15n=75/0.15 n=500